the characteristic of the equation PQ=x and the integral surface which passes through the curve y=1/2 z=x are =?


1
Expert's answer
2021-11-24T14:56:16-0500
"f(x, y, z, p, q)=pq-x=0""f_x=-1, f_y=0, f_z=0, f_p=q, f_q=p"




"\dfrac{dx}{f_p}=\dfrac{dy}{f_q}=\dfrac{dz}{pf_p+qf_q}=\dfrac{dp}{-(f_x+pf_z)}=\dfrac{dq}{-(f_y+qf_z)}"



"\dfrac{dx}{q}=\dfrac{dy}{p}=\dfrac{dz}{2pq}=\dfrac{dp}{1}=\dfrac{dq}{0}"


"dx\/q=dp"

"dq=0\implies q=c"

"p=x\/c+c_1"

"z=x^2\/(2c)+c_1x+c_3"


"dy=pdp"

"p^2=2y+c_2"

"(x\/c+c_1)^2=2y+c_2"


"F(z-x^2\/(2c)-c_1x,(x\/c+c_1)^2-2y)=0"


we have y=1/2 and z=x, so


"F(x(1-c_1)-x^2\/(2c),(x\/c+c_1)^2-1)=0"